Canon by 12 Parts (Variant in Major) on 1 Samuel 7,3

[figure: {PTL o.m.1 \atop 5.4.1989}; Sa-; mu-; el; spoke; and; said]
[figure: to; all; the; house; of; Is-; ra-; el:]

[figure: If; you; are; re-; turn-; ing; sure-; ly]
[figure: to; the; LORD; with; all; your; hearts,]

[figure: then; put; a-; way]
[figure: the; for-; eign; i-; dols,]
[figure: and; the; Ash-; ta-; roth,]
[figure: and; ded-; i-; cate; your-; selves]
[figure: to; the; LORD,]
[figure: and; serve; on-; ly; him,]
[figure: and; tru-; ly; he; will; save; you]
[figure: Accompaniment:; from; D; the; _{fis_{7}}; Phi-; b; G; lis-; b; tine; A; pow-; D; er.; D]